OpenStudy (anonymous):

How can I graph 5x+2y=8 .

OpenStudy (anonymous):

It's easier to put it into slope intercept form, which would be y= -2.5x+4 just enter that into your calculator, or if graphing on paper, start at (0,4) and follow rise over run.

OpenStudy (nincompoop):

You don't need to convert it into slope-intercept form. you can directly solve for x and y. separately, solve for x by making y equal to zero then solve for y by making x zero. you end up with the first point (x, 0) and second point (0, y) The first suggested way of approaching this problem works as well, but it just gives you more steps to work on.

OpenStudy (anonymous):

What nin said

OpenStudy (anonymous):

@nincompoop : Can show me an example please?

OpenStudy (nincompoop):

let us say we have an equation of a line in standard form (this is what standard form looks like: Ax+By = C) 10x + 4y = 16 As I said earlier solve for x and y separately. When solving for x make sure to make y equal to zero 10x + B(0) = 16 10x = 16 divide both sides by 2 to get x x = 8 this means that when y = 0 you get x = 8. Make this as your first point (8, 0) Then solve for y. When solving for y, make sure you make x equal to zero. 10(0) +4y = 16 4y = 16 divide both sides by 4 to get y y = 4 this means that when x = 0 you get y = 4. Make this as your second point (0, 4) plot these two points in a cartesian graph and draw a straight line connecting the two label the points and put the equation where everyone can see that the line has the equation 10x + 4y = 16

OpenStudy (nincompoop):

it seems a lot of work, but that is because you're not used to it yet. This thing becomes easy you would be able do it in your mind and only need to write down very little but important information.

OpenStudy (nincompoop):

thank you for your patience and I hope that you learned something new.

OpenStudy (anonymous):

I see now, And I actually understood this. It helped a lot, Thank you! (: But theres also one thing on solving for x , When I do 5x+2(0)=8 . I get 5x=8 , Now how would I simplify that down?

OpenStudy (nincompoop):

yes you always need to simplify down to just x or y 5x means you need to divide both sides by 5 to make it just x

OpenStudy (nincompoop):

as you can see, I made sure that I had just x and y

OpenStudy (anonymous):

But when I divide by 5 on both sides , I can divide the 5x and get x . But when I divide 5 by 8 , What would I get?

OpenStudy (nincompoop):

8/5

OpenStudy (nincompoop):

if you want to convert that into decimal or mixed fraction go ahead, but you don't really have to.

OpenStudy (anonymous):

8/5 it's irreducible, it's like 1.6 or something lol, you can just use normal division to figure it out if you're uncertain what it actually is.